A skier with a mass of 62 π‘˜π‘” is sliding down a snowy slope. Find the coefficient of kinetic friction for the skier if friction is known to be 45.0 𝑁.


1
Expert's answer
2021-10-02T10:47:04-0400

If the speed of descending is constant, the coefficient of kinetic friction is


"f=\mu mg\cos\theta,\\\space\\
f=mg\sin\theta.\\\space\\
\theta=\arcsin\frac{f}{mg}=4.2°.\\\space\\
\mu=\frac{f}{mg\cos\theta}=0.07."